# Chaos in hyperscaling violating Lifshitz theories

### v1

@@ Abstract
We holographically study quantum chaos in hyperscaling-violating Lifshitz (HVL) theories (with charge). Particularly, we present a detailed computation of the out-of-time ordered correlator (OTOC) via the shockwave analysis in the bulk HVL geometry with a planar horizon topology. We also compute the butterfly velocity ($v_{B}$) using the entanglement wedge reconstruction and find that the result matches the one obtained from shockwave analysis. Furthermore, we analyze in detail, the behavior of $v_{B}$ with respect to the dynamical critical exponent (z), hyperscaling-violating parameter ($\theta$), charge (Q) and the horizon radius ($r_{h}$). We interestingly find non-monotonic behavior of $v_{B}$ with respect to z (in the allowed region and for certain (not all) fixed, permissible values of $\theta$, Q and $r_{h}$) and $\theta$ (in the allowed region and for certain (not all) fixed, permissible values of z, Q and $r_{h}$). Moreover, $v_{B}$ is found to monotonically decrease with an increase in charge (for all permissible, fixed values of z, $\theta$ and $r_{h}$), whereas it is found to monotonically increase with $r_{h}$ (for all fixed, permissible values of z, $\theta$ and Q). Unpacking these features can offer some valuable insights into the chaotic nature of HVL theories.
@@ I   Introduction
Classically, chaos is characterized by an exponential divergence between two phase space trajectories, with the Lyapunov exponent $\lambda_{L}$ parametrizing the rate of this divergence. However, characterization of quantum chaos is a lot more challenging, because the sensitivity in initial conditions cannot be directly computed due to the uncertainty principle. Initially, characterization of quantum chaos was based on comparison between the system’s spectrum of energies and the spectrum of random matrices [3]. However, a new approach was developed in [4] (in the context of semi-classical systems) and [5]. Quantum chaos can be characterized by the strength of the commutator between two generic operators V and W separated by time t. To be precise, we consider the following quantity: $-[W(t),V(0)]^{2}\right>_{\beta}$, where expectation value is taken in the thermal state $\beta$. One can also consider the spreading of chaos in spatial directions, by taking the two operators to be also spatially separated. In such a case, the commutator is given by (12), where $v_{B}$ is the butterfly velocity. Roughly speaking, $v_{B}$ characterizes the rate at which the information about the applied perturbation ”spreads” (or scrambles among the local degrees of freedom) in the system (this notion is made precise in III). Focus is laid on $v_{B}$ in this work.
@@
Gauge/gravity duality ([6] [7]) is the idea that a gravitational theory in a (d+1) dimensional bulk spacetime is dual to a d dimensional quantum theory (without gravity) residing at the asymptotic boundary of the bulk spacetime. AdS/CFT correspondence ([8], [9] [10]) is the most well understood realization of this duality. In the framework of AdS/CFT correspondence, a gravitational theory in asymptotically AdS spacetime in the bulk is dual to a conformal field theory (CFT). However, the general principles of gauge/gravity duality can be extended to bulk spacetimes that are not asymptotically AdS and equivalently boudary theories that are not conformally invariant. Lifshitz theories are an example of such theories that are not conformally invariant. These theories exhibit an anisotropic scaling symmetry between space and time, parametrized by the dynamical critical exponent z, i.e. $\{t,x\}\rightarrow\{\alpha^{z}t,\alpha x\}$. Theories with $z\neq 1$ do not support Lorentz invariance and instead possess non-relativistic symmetries. Also, for $z\neq 1$, these theories describe quantum critical systems (non-relativistic). Bulk gravitational theory dual to Lifshitz theory was first proposed in [11]. (see also [12] [13] [14][15], [16]). The anisotropic scaling symmetry of the boundary theory can be geometrically realized in the bulk via these so-called Lifshitz spacetimes.
@@
Quantum critical point may be characterized by several other critical exponents that satisfy certain relations and thereby are dependent on each other. Hyperscaling relations are a certain class of critical exponent relations in which there is an explicit appearance of dimensionality of the space [17]. There are critical theories that violate this hyperscaling relation. For such theories the heat capacity $C_{v}\sim T^{(d-1-\theta)/z}$ ([18],[19]), where $\theta$ is called the hyperscaling-violating parameter. Essentially, $\theta$ reduces the effective dimensionality of the theory. For hyperscaling-violating and Lifshitz (HVL) theoreies, the dual bulk black hole solutions have been obtained in several works [20] [21] [22] [23] [24] [25]. In this paper, we closely follow the discussion in [25].
@@
In recent years, the study of chaotic dynamics of strongly-coupled many-body quantum systems via gauge-gravity duality has witnessed a significant interest [26][27][28] [29]. For a comprehensive review of holographic chaos, refer [30]. Through [31], it was established that black holes are the fastest scramblers in nature. Thus, in the context of AdS/CFT, the rapid thermalization of a local perturbation in the boundary CFT is understood through the fast scrambling dynamics of a black hole in the bulk. It is also known that black holes in holography can be characterized via quantum chaos [32][31]. Possessing maximal chaos is thought of as a criterion for a QFT to have a gravity dual. [8] [26][33].
@@
In this work, we holographically study quantum chaos in HVL theories (with charge). In particular, we compute the out-of-time-ordered correlator (OTOC) (we make the notion precise in III) via shockwave analysis in the bulk geometry (for planar horizon topology) and thereby compute the various chaos parameters, namely the Lyapunov exponent ($\lambda_{L}$), butterfly velocity ($v_{B}$) and the scrambling time ($t_{*}$) for the HVL theories. One of the motivations for studying OTOCs in HVL theories is the recent proposals for experimentally measuring OTOCs in quantum systems [34] [35] [36] [37]. We also compute $v_{B}$ using the entanglement wedge reconstruction and find the results obtained from both the methods, to match. However, the main objective of this work is to point out the interesting features in the behavior of $v_{B}$ when varied with respect to z, $\theta$, Q and $r_{h}$. Interestingly, we find certain non-monotonicities in $v_{B}$ when varied against z and $\theta$.
Recently, analysis of OTOC, entanglement wedge and pole skipping for Lifshitz geometries was presented in [38]. We find our results to be consistent with theirs, for $\theta=0$.
@@
The structure of this paper is as follows: In II, we discuss the charged hyperscaling-violating Lifshitz black hole solution and also discuss the constraints on z and $\theta$ that arise from the null-energy condition. In III, we provide a detailed analysis of computing the OTOC for HVL theories via the shockwave analysis. In IV, we compute $v_{B}$ using the entanglement wedge reconstruction and find the results of III and IV to match. In V, we discuss the variation of $v_{B}$ with respect to the various parameters, namely z, $\theta$, Q and $r_{h}$. We conclude in VI.
@@ II   Background
k is the factor that characterizes the topology of the horizon. $k=\{-1,0,1\}$ for hyperbolic, planar and spherical horizon topology respectively. $r_{F}$ is the large radius upto which the black hole geometry is considered to be valid. $r_{F}$ corresponds to the UV cutoff scale in the boundary. L is the bulk curvature radius. The temperature can be computed from the blackening function.
@@
Here F, H and K are the field strengths corresponding to the three Abelian gauge fields A,B,C, given as $F=dA$, $H=dB$ and $K=dC$. $\phi$ is a real scalar field known as the dilaton field. X, Y and Z are the respective coupling functions of the gaugle fields and the dilaton field. Note that the gauge field A supports the Lifshitz asymptotics, B supports the other non-planar horizon topologies (hyperbolic and spherical) and C allows for a non-zero electric charge. In this work we deal only with planar horizon topology ($k=0$) and set $L=1$. For $k=0$ and $L=1$, f(r) and temperature are given as
@@ II.1   Constraints on z and $\theta$
These constraints play an important role in the discussion in V.
@@ III   Out-of-time ordered correlator (OTOC)
For two spatially separated generic Hermitian operators V and W, quantum chaos can be characterized by the following commutator $[W(0,-t),V(x,0)]$. Note that, since we are interested in thermal systems, we take an expectation value of the commutator in a thermal state $\beta$. However, the thermal expectation value may attain a negative sign which can thereby lead to cancellations. To overcome this, we take the square of the commutator and thus the precise quantity we consider, is the following.
@@ III.1   Kruskal extension and the shockwave analysis
In this subsection we compute the various chaos parameters (and thereby the OTOC) for HVL theories by incorporating the shockwave analysis in the bulk. We work with the two sided eternal black hole geometry, which corresponds to a thermo-field double state (denoted as |$\ket{TFD}$) in the boundary. Note that we will consider planar horizon topology (i.e. $k=0$) for our analysis throughout.
In order to work in the two-sided geometry, we write the metric (1) in the Kruskal co-ordinates. The Kruskal co-ordinate transformation is given as
@@
$r_{*}\approx\frac{1}{f^{\prime}(r_{h})r_{h}^{z+1}}ln(r-r_{h})$ (21)
@@
$u\,(v=0)$$v\,(u=0)$$h(x)$
Figure 1: The Penrose diagram, explaining the effect of the shockwave created near the horizon. The red line represents the shockwave created by the perturbation particle W, which is inserted at a sufficiently earlier time. In the absence of the shockwave, particle V would have emerged from the past interior and would have reached the boundary at $t=0$. However, due to the shockwave, the V particle suffers a shift in it’s trajectory, which is parametrized by h(x). This causes a delay in the appearance of operator V at the boundary. If we apply the W perturbation early enough (i.e. $|t|\gtrsim|t_{*}|$), the trajectory of the particle gets shifted to the point that it gets engulfed by the future interior and the V particle fails to reach the boundary, altogether.
Here, the step function $\Theta(u)$ accounts for the fact that the causal future of the particle is affected after it encounters the horizon $u=0$ (and thus there’s no affect on the causal past). h(x,t) is a function that characterizes the amount of shift and it is to be determined by the Einstein equation of motion. To understand the relation between this function h(x,t) and the commutator (and thereby the OTOC), we consider the bulk picture of the two states |$\ket{\psi_{1}}$ and |$\ket{\psi_{2}}$. The bulk picture of |$\ket{\psi_{1}}$ is as follows: Initially,the system is prepared in such a way that the V particle from the past interior reaches the boundary at $t=0$. However, the insertion of the operator W at a sufficiently earlier time makes the particle encounter a localized shock at the horizon, which results in the shift of particle’s trajectory, which in turn, causes a time delay in the appearance of the operator V at the boundary. On the other hand, the state |$\ket{\psi_{2}}$ corresponds to the V particle reaching the boundary at $t=0$, ”after” suffering the shift due to the shock. Thus, the overlap between the two states (and thus the OTOC) keeps decreasing for earlier and earlier times. After a certain past time, the shift suffered by the particle is such that the particle gets trapped inside the future interior and then it cannot escape to the boundary, making the overlap and thus the OTOC completely vanish. This time is given by the scrambling time, $|t|\gtrsim|t_{*}|$.
@@
$A(0)=\lim_{u\to 0}A=\frac{r_{h}^{-1-2d}\left(\frac{r_{h}}{r_{F}}\right)^{-\frac{2\theta}{-1+d}}\left(-r_{h}^{2(d+z)}(-1+d+z-\theta)-q^{2}r_{h}^{4+2\theta}(3-d-z+\theta)\right)}{2\alpha^{2}}$ (35)
@@
$\partial_{u}\partial_{v}B(0)=\frac{-2r_{h}\left(\frac{r_{h}}{r_{F}}\right)^{-\frac{2\theta}{-1+d}}+2r_{h}^{2}\left(\frac{r_{h}}{r_{F}}\right)^{-1-\frac{2\theta}{-1+d}}\theta}{(-1+d)\,r_{F}}$ (37)
@@
$\begin{split}t_{*}=&\,\frac{1}{2}r_{h}^{1+z}\left(-2q^{2}r_{h}^{3-2d-2z+2\theta}(-2+d+z-\theta)+(1+q^{2})\left(\frac{1}{r_{h}}\right)^{d+z-\theta}(-1+d+z-\theta)\right)\\ &\times\log\left(\frac{\left(-r_{h}^{2(d+z)}(-1+d+z-\theta)-q^{2}r_{h}^{4+2\theta}(3-d-z+\theta)\right)}{4G_{N}\pi E_{o}r_{h}^{5}\left(-2q^{2}r_{h}^{3-d-z+2\theta}(-2+d+z-\theta)+(1+q^{2})r_{h}^{\theta}(-1+d+z-\theta)\right)^{2}}\right)\end{split}$ (38)
@@
$t_{*}=\frac{1}{2}\left(-1+d+z-\theta+q^{2}(3-d-z+\theta)\right)\log\left(\frac{1}{4G_{N}\pi E_{o}\left(1-d-z+\theta-q^{2}(3-d-z+\theta)\right)}\right)$ (41)
@@
Note that, none of the chaos parameters depend on the cutoff scale $r_{F}$.

### v2

@@ Abstract
We holographically study quantum chaos in hyperscaling-violating Lifshitz (HVL) theories (with charge). Specifically, we present a detailed computation of the out-of-time ordered correlator (OTOC) via shockwave analysis in the bulk HVL geometry with a planar horizon topology. We also compute the butterfly velocity ($v_{B}$) using the entanglement wedge reconstruction and find that it matches the result obtained from the shockwave analysis. Using a recently developed thermodynamic dictionary for HVL theories, we express $v_{B}$ purely in terms of boundary thermodynamic variables. Furthermore, we analyze in detail the behavior of $v_{B}$ with respect to the dynamical critical exponent ($z$), hyperscaling-violating parameter ($\theta$), entropy (more precisely, the ratio of entropy to the central charge, $\tilde{S}$), and charge (more precisely, the ratio of charge to the central charge, $\tilde{Q}$). Interestingly, $v_{B}$ varies non-monotonically with $z$ for $\tilde{S}<1$, whereas it increases monotonically with $z$ for $\tilde{S}\geq 1$. Additionally, $v_{B}$ varies non-monotonically with $\theta$ for non-zero charge. Moreover, $v_{B}$ monotonically increases with $\tilde{S}$ and decreases with $\tilde{Q}$ for all allowed values of $z$ and $\theta$. All these features are reported for combinations {$z$, $\theta$, $\tilde{S}$, $\tilde{Q}$} for which the temperature is positive, the null energy condition is satisfied, and $v_{B}$ is not superluminal. Unpacking the non-monotonicities in $v_{B}$ can offer interesting insights into these theories.
@@ I   Introduction
Classically, chaos is characterized by an exponential divergence between two phase space trajectories, with the Lyapunov exponent $\lambda_{L}$ parametrizing the rate of this divergence. However, characterization of quantum chaos is a lot more challenging, because the sensitivity in initial conditions cannot be directly computed due to the uncertainty principle. Initially, characterization of quantum chaos was based on comparison between the system’s spectrum of energies and the spectrum of random matrices [3]. However, a new approach was developed in [4] (in the context of semi-classical systems) and [5]. Quantum chaos can be characterized by the strength of the commutator between two generic operators V and W separated by time t. To be precise, the following quantity is considered: $-[W(t),V(0)]^{2}\right>_{\beta}$, where expectation value is taken in the thermal state $\beta$. One can also consider the spreading of chaos in spatial directions, by taking the two operators to be also spatially separated. In such a case, the commutator is given by (15), where $v_{B}$ is the butterfly velocity. Roughly speaking, $v_{B}$ characterizes the rate at which the information about the applied perturbation ”spreads” (or scrambles among the local degrees of freedom) in the system (this notion is made precise in III). Focus is laid on $v_{B}$ in this work.
@@
Gauge/gravity duality ([6], [7]) is the idea that a gravitational theory in a (d+1) dimensional bulk spacetime is dual to a d dimensional quantum theory (without gravity) residing at the asymptotic boundary of the bulk spacetime. AdS/CFT correspondence [8, 9, 10] is the most well understood realization of this duality. In the framework of AdS/CFT correspondence, a gravitational theory in asymptotically AdS spacetime in the bulk is dual to a conformal field theory (CFT). However, the general principles of gauge/gravity duality can be extended to bulk spacetimes that are not asymptotically AdS and equivalently boudary theories that are not conformally invariant. Lifshitz theories are an example of such theories that are not conformally invariant. These theories exhibit an anisotropic scaling symmetry between space and time, parametrized by the dynamical critical exponent z, i.e. $\{t,x\}\rightarrow\{\alpha^{z}t,\alpha x\}$. Theories with $z\neq 1$ do not support Lorentz invariance and instead possess non-relativistic symmetries. Also, for $z\neq 1$, these theories describe quantum critical systems (non-relativistic). Bulk gravitational theory dual to Lifshitz theory was first proposed in [11]. (see also [12, 13, 14, 15, 16]). The anisotropic scaling symmetry of the boundary theory can be geometrically realized in the bulk via these so-called Lifshitz spacetimes.
@@
Quantum critical point may be characterized by several other critical exponents that satisfy certain relations and thereby are dependent on each other. Hyperscaling relations are a certain class of critical exponent relations in which there is an explicit appearance of dimensionality of the space [17]. There are critical theories that violate this hyperscaling relation. For such theories the heat capacity $C_{v}\sim T^{(d-1-\theta)/z}$ ([18],[19]), where $\theta$ is called the hyperscaling-violating parameter. Essentially, $\theta$ reduces the effective dimensionality of the theory. For hyperscaling-violating and Lifshitz (HVL) theoreies, the dual bulk black hole solutions have been obtained in several works [20, 21, 22, 23, 24, 25]. In this paper, we closely follow the discussion in [25].
@@
In recent years, the study of chaotic dynamics of strongly-coupled many-body quantum systems via gauge-gravity duality has witnessed a significant interest [26, 27, 28, 29]. For a comprehensive review of holographic chaos, refer [30]. Through [31], it was established that black holes are the fastest scramblers in nature. Thus, in the context of AdS/CFT, the rapid thermalization of a local perturbation in the boundary CFT is understood through the fast scrambling dynamics of a black hole in the bulk. It is also known that black holes in holography can be characterized via quantum chaos [31], [32]. Possessing maximal chaos is thought of as a criterion for a QFT to have a gravity dual [8], [26], [33].
@@
In this work, we holographically study quantum chaos in HVL theories (with charge). In particular, we compute the out-of-time-ordered correlator (OTOC) (we make the notion precise in III) via shockwave analysis in the bulk geometry (for planar horizon topology) and thereby compute the various chaos parameters, namely the Lyapunov exponent ($\lambda_{L}$), butterfly velocity ($v_{B}$) and the scrambling time ($t_{*}$) for the HVL theories. One of the motivations for studying OTOCs in HVL theories is the recent proposals for experimentally measuring OTOCs in quantum systems [34, 35, 36, 37]. We also compute $v_{B}$ using the entanglement wedge reconstruction and find the results obtained from both the methods, to match. Furthermore, we study how boundary thermodynamics and theory parameters (z and $\theta$) affect information scrambling. To achieve this, we use a recently developed thermodynamic dictionary for the HVL theories [38] to express $v_{B}$ in terms of boundary thermodynamic variables. We study the variation of $v_{B}$ with respect to the theory parameters $z$ and $\theta$, as well as the boundary thermodynamic variables: entropy over central charge ($\tilde{S}\sim S/C$) and charge over central charge ($\tilde{Q}\sim Q/C$). Several interesting features are obtained.
@@
The structure of this paper is as follows: In II, we discuss the charged hyperscaling-violating Lifshitz black hole solution and also discuss the constraints on z and $\theta$ that arise from the null-energy condition. In III, we provide a detailed analysis of computing the OTOC for HVL theories via the shockwave analysis. In IV, we compute $v_{B}$ using the entanglement wedge reconstruction and find the results of III and IV to match. In V, we discuss the thermodynamic dictionary and express $v_{B}$ in terms of the boundary thermodynamic variables. In VI, we discuss the variation of $v_{B}$ with respect to the various z, $\theta$, $\tilde{S}$, and $\tilde{Q}$. We conclude in VII.
@@ II   Background
$r_{h}$ is the horizon radius, d is the number of boundary dimensions, q is the charge parameter, k is the factor that characterizes the topology of the horizon ($k=\{-1,0,1\}$ for hyperbolic, planar and spherical horizon topology respectively), $r_{F}$ is the large radius upto which the black hole geometry is considered to be valid, ($r_{F}$ corresponds to the UV cutoff scale in the boundary), and L is the bulk curvature radius. The temperature can be computed from the blackening function.
@@
Here F, H and K are the field strengths corresponding to the three Abelian gauge fields A,B,C, given as $F=dA$, $H=dB$ and $K=dC$. $\phi$ is a real scalar field known as the dilaton field. X, Y and Z are the respective coupling functions of the gaugle fields and the dilaton field given as
$X=X_{o}e^{\lambda_{1}\phi},\hskip 8.5359ptY=Y_{o}e^{\lambda_{2}\phi},\hskip 8.5359ptZ=Z_{o}e^{\lambda_{3}\phi}$ (5)
Here, $X_{o}$, $Y_{o}$, and $Z_{o}$ are positive parameters that quantify the strength of coupling between gravity and the three gauge fields respectively, and $\lambda_{i}s$ are constants given as
$\lambda_{1}=\frac{-2(d-1-\theta+\theta/(d-1))}{\tilde{\gamma}},\hskip 8.5359pt\lambda_{2}=\frac{-2(d-2)(d-\theta-1)}{\tilde{\gamma}(d-1)},\hskip 8.5359pt\lambda_{3}=\frac{\tilde{\gamma}}{d-\theta-1}$ (6)
where $\tilde{\gamma}$ is given as
$\tilde{\gamma}=\sqrt{2(d-\theta-1)\left(z-1-\frac{\theta}{d-1}\right)}$ (7)
Note that the gauge field A supports the Lifshitz asymptotics, B supports the other non-planar horizon topologies (hyperbolic and spherical) and C allows for a non-zero electric charge. In this work we deal only with planar horizon topology ($k=0$) and set $L=1$. For $k=0$ and $L=1$, f(r) and temperature are given as
@@ II.1   Constraints on z and $\theta$
These constraints play an important role in the discussion in VI.
@@ III   Out-of-time ordered correlator (OTOC)
For two spatially separated generic Hermitian operators V and W, quantum chaos can be characterized by the following commutator $[W(0,-t),V(x,0)]$. Note that, since it is of interest to study thermal systems, an expectation value of the commutator is taken in a thermal state $\beta$. However, the thermal expectation value may attain a negative sign which can thereby lead to cancellations. To overcome this, the commutator is squared and thus the precise quantity to be considered, is the following.
@@ III.1   Kruskal extension and the shockwave analysis
In this subsection, we compute the various chaos parameters for HVL theories via the shockwave analysis in the bulk. It is standard to work with the two-sided eternal black hole geometry, which corresponds to a thermo-field double state (denoted as |$\ket{TFD}$) in the boundary. Note that we will consider planar horizon topology (i.e. $k=0$) for our analysis throughout.
In order to work in the two-sided geometry, the metric (1) is written in the Kruskal co-ordinates. The Kruskal co-ordinate transformation is given as
@@
$r_{*}\approx\frac{1}{f^{\prime}(r_{h})r_{h}^{z+1}}\log(r-r_{h})$ (24)
@@
$u\,(v=0)$$v\,(u=0)$$h(x)$
Figure 1: The Penrose diagram, explaining the effect of the shockwave created near the horizon. The red line represents the shockwave created by the perturbation particle W, which is inserted at a sufficiently earlier time. In the absence of the shockwave, particle V would have emerged from the past interior and would have reached the boundary at $t=0$. However, due to the shockwave, the V particle suffers a shift in its trajectory, which is parametrized by $h(x)$. This causes a delay in the appearance of operator V at the boundary. If the W perturbation is applied early enough (i.e., $|t|\gtrsim|t_{*}|$), the trajectory of the particle gets shifted to the point that it gets engulfed by the future interior and the V particle fails to reach the boundary altogether.
@@
Here, the step function $\Theta(u)$ accounts for the fact that the causal future of the particle is affected after it encounters the horizon $u=0$ (and thus there’s no affect on the causal past). h(x,t) is a function that characterizes the amount of shift and it is to be determined by the Einstein equation. To understand the relation between this function h(x,t) and the commutator (and thereby the OTOC), consider the bulk picture of the two states |$\ket{\psi_{1}}$ and |$\ket{\psi_{2}}$. The bulk picture of |$\ket{\psi_{1}}$ is as follows: Initially,the system is prepared in such a way that the V particle from the past interior reaches the boundary at $t=0$. However, the insertion of the operator W at a sufficiently earlier time makes the particle encounter a localized shock at the horizon, which results in the shift of particle’s trajectory, which in turn, causes a time delay in the appearance of the operator V at the boundary. On the other hand, the state |$\ket{\psi_{2}}$ corresponds to the V particle reaching the boundary at $t=0$, ”after” suffering the shift due to the shock. Thus, the overlap between the two states (and thus the OTOC) keeps decreasing for earlier and earlier times. After a certain past time, the shift suffered by the particle is such that the particle gets trapped inside the future interior and then it cannot escape to the boundary, making the overlap and thus the OTOC completely vanish. This time is given by the scrambling time, $|t|\gtrsim|t_{*}|$.
@@
$A(0)=\lim_{u\to 0}A=\frac{r_{h}^{-1-2d}\left(\frac{r_{h}}{r_{F}}\right)^{-\frac{2\theta}{d-1}}\left(-r_{h}^{2(d+z)}(d-1+z-\theta)-q^{2}r_{h}^{4+2\theta}(3-d-z+\theta)\right)}{2\alpha^{2}}$ (38)
@@
$\partial_{u}\partial_{v}B(0)=\frac{-2r_{h}\left(\frac{r_{h}}{r_{F}}\right)^{-\frac{2\theta}{d-1}}+2r_{h}^{2}\left(\frac{r_{h}}{r_{F}}\right)^{-1-\frac{2\theta}{d-1}}\theta}{(d-1)\,r_{F}}$ (40)
@@
$\begin{split}t_{*}=&\,\frac{1}{2}r_{h}^{1+z}\left(-2q^{2}r_{h}^{3-2d-2z+2\theta}(d-2+z-\theta)+(1+q^{2})\left(\frac{1}{r_{h}}\right)^{d+z-\theta}(d-1+z-\theta)\right)\\ &\times\log\left(\frac{\left(r_{h}^{2(d+z)}(1-d-z+\theta)+q^{2}r_{h}^{4+2\theta}(d-3+z-\theta)\right)}{4G_{N}\pi E_{o}r_{h}^{5}\left(-2q^{2}r_{h}^{3-d-z+2\theta}(d-2+z-\theta)+(1+q^{2})r_{h}^{\theta}(d-1+z-\theta)\right)^{2}}\right)\end{split}$ (41)
@@
$t_{*}=\frac{1}{2}\left(d-1+z-\theta+q^{2}(3-d-z+\theta)\right)\log\left(\frac{1}{4G_{N}\pi E_{o}\left(1-d-z+\theta-q^{2}(3-d-z+\theta)\right)}\right)$ (44)
@@
Note that for $\theta=0$, $r_{h}=1$, and $q=0$, the result (42) matches the result obtained in [39]. Also, as an important sidenote, we point out that the butterfly velocity (42) is equal to the velocity of entanglement growth during thermalization at the saturation point  [^1] Author would like to thank Mohammad Reza Mohammadi Mozaffar for pointing this out. (for $q=0$) [40] , suggesting an interesting connection between the two notions.